OpenStudy (anonymous):

Marks: 2 Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66. Choose one answer. a. 70 b. 72 c. 111 d. 114

OpenStudy (lgbasallote):

find the common difference first \[A_{27} = A_1 + (n - 1)d\] \[66 = -12 + (28 - 1)d\]

OpenStudy (anonymous):

ok then what next

OpenStudy (lgbasallote):

find d

OpenStudy (anonymous):

how do I do that

OpenStudy (lgbasallote):

\[66 = 12+ 27d\] isolate d

OpenStudy (anonymous):

I tried but i never get a whole number

OpenStudy (anonymous):

someone please help

OpenStudy (lgbasallote):

what did you get?

OpenStudy (anonymous):

2.8

OpenStudy (lgbasallote):

\[66 - 12 = 27d\] \[54 = 27d\]

OpenStudy (lgbasallote):

try that...that becomes whole

OpenStudy (anonymous):

but it was already a -12 so don't you add them

OpenStudy (lgbasallote):

it was originally \[66 = 12 + 27d\] i subtracted 12 from both sides \[66 - 12 = 27d\]

OpenStudy (anonymous):

where did you get the 12 from?

OpenStudy (lgbasallote):

ugh...i messed up my equation =_= should be \[66 = -12 + (27 - 1)d\] because there are 27 terms *facepalm* sorry

OpenStudy (anonymous):

Thats ok that what i am confused on thats how I got 2.8

OpenStudy (lgbasallote):

\[66 = -12 + 26d\] \[66 + 12 = 26d\] you should get a whole number now

OpenStudy (anonymous):

ok I got 3

OpenStudy (lgbasallote):

so common denominator is 3... now we solve for 42nd term \[A_{42} = A_1 + (n - 1)d\]

OpenStudy (lgbasallote):

a1 would be -12 n would be 42 d would be 3

OpenStudy (anonymous):

so it would be 111

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thanks!